# Improving Quotation Attribution with Fictional Character Embeddings

### v1

@@ Abstract
Humans naturally attribute utterances of direct speech to their speaker in literary works. When attributing quotes, we process contextual information but also access mental representations of characters that we build and revise throughout the narrative. Recent methods to automatically attribute such utterances have explored simulating human logic with deterministic rules or learning new implicit rules with neural networks when processing contextual information. However, these systems inherently lack character representations, which often leads to errors on more challenging examples of attribution: anaphoric and implicit quotes. In this work, we propose to augment a popular quotation attribution system, BookNLP, with character embeddings that encode global information of characters. To build these embeddings, we create DramaCV, a corpus of English drama plays from the 15th to 20th century focused on Character Verification (CV), a task similar to Authorship Verification (AV), that aims at analyzing fictional characters. We train a model similar to the recently proposed AV model, Universal Authorship Representation (UAR), on this dataset, showing that it outperforms concurrent methods of characters embeddings on the CV task and generalizes better to literary novels. Then, through an extensive evaluation on 22 novels, we show that combining BookNLP’s contextual information with our proposed global character embeddings improves the identification of speakers for anaphoric and implicit quotes, reaching state-of-the-art performance. Code and data will be made publicly available.
@@ 1 Introduction
Recent approaches to quotation attribution typically succeed at attributing explicit quotes but struggle when it comes to anaphoric and implicit quotes Muzny et al. (2017b); Cuesta-Lazaro et al. (2022); Vishnubhotla et al. (2023); Su et al. (2023). Explicit utterances occur when the narrator indicates the speaker of a quote with a speech verb and a named mention, while anaphoric quotes are introduced with a speech verb and a pronoun or common noun. When no narrative information is given on the speaker of the quote, we refer to those as implicit quotes. Examples of such quotes are given in Figure 1.
@@
Attributing anaphoric and implicit quotes usually requires fine-grained understanding of contextual information such as linking pronouns to their canonical entity or discerning discussion patterns. Accessing information of fictional characters (style, persona, gender) also help disambiguate challenging examples of attribution. Variations of writing style and long-range contextual dependencies, ubiquitous in fictional works, add an extra layer of complexity, making attribution of non-explicit quotes a complex task to solve even with recent tools. Sims and Bamman (2020); Vishnubhotla et al. (2022)
@@
While some of these shortcomings have been addressed by improving link prediction between pronouns/common nouns and canonical characters Vishnubhotla et al. (2023) or by exploiting better contextual representations Cuesta-Lazaro et al. (2022); Su et al. (2023), we instead show that combining representations encoding global stylistic information and topical preferences of fictional characters with contextual information improves the identification of speakers of anaphoric and implicit quotes. In particular, we combine these embeddings with BookNLP [^1] https://github.com/booknlp/booknlp, a widely used NLP pipeline designed for literary texts. As explicit utterances can easily be attributed to their speaker, we assume those quotes given, and use them to inject global character-level information into BookNLP.
@@
Among methods to represent characters from the words they utter, previous works Li et al. (2023); Aggazzotti et al. (2023); Michel et al. (2024) have explored embeddings encoding topical preferences using SBERT Reimers and Gurevych (2019) or stylistic information using Universal Authorship Representations (UAR) Rivera-Soto et al. (2021), a fine-tuned variant of SBERT trained on the Authorship Verification (AV) task with millions of Reddit users. However, considering the discrepancies between Reddit and literary discourse, we question the effectiveness of these embeddings to represent characters. To overcome this potential limitation, we build a variant of UAR tailored for literary texts and evaluate those representation models on quotation attribution and Character Verification (CV), a task similar to AV, aiming at distinguishing characters in literary works using their utterances within authorial bounds.
Our proposed models significantly improve CV on drama plays and novels, and reaches state-of-the-art quotation attribution accuracy of non-explicit quotes when used in combination with BookNLP. Besides, we show that using models tailored for literary text leads to better quotation attribution performance. To sum up, our contributions are:
@@
  • We construct DramaCV, a corpus of English drama plays, dating from the 15th to 20th century focused on CV, and train variants of UAR on this dataset, showing that our proposed models significantly improves CV on DramaCV and also generalize better to novels.
@@
  • Through an extensive evaluation, we show that BookNLP can be easily extended to improve its current performance on the Project Dialogism Novel Corpus (PDNC) Vishnubhotla et al. (2022), and reaches state-of-the-art accuracy on non-explicit quotes when trained in combination with global character representations computed from their explicit quotes.
@@ 2 Related Work
Fictional Character Embeddings.   Understanding fictional characters objectives, personas and style often comes naturally when reading novels. When accessing information of the fictional world and interactions between characters provided throughout the narrative, we infer mental representations of characters that are revised when accessing new information Gernsbacher et al. (1992); Culpeper (1996). Previous computational works on this aspect of fictional character understanding have explored representing fictional characters in a latent space, often targeting a specific aspect, such as persona, style or descriptive elements. These computational representations of characters are often used for further analysis on large literary corpora, trying to understand character distinctiveness between authors Bamman et al. (2014) or narrative comprehension Brahman et al. (2021). Other works focus on using fictional characters embedings to improve character-level tasks such as coreference resolution and character linking Li et al. (2023) or contextual and book-level tasks such as narrative question answering Inoue et al. (2022).
Here, we propose to represent character from what characters say and how they say it, building literary embeddings encoding style and topical preferences of characters, derived from models trained on Authorship Verification or Character Verification. We define Character Verification (CV) as a task closely related to Authorship Verification, aiming at distinguishing characters in literary works, within authorial bounds. Most works focusing on CV analyze drama characters because of the availability of large annotated corpora and aim at capturing syntactic, lexical and phonological variations that occur in their direct speech using statistical measures of distinctiveness Dinu and Uban (2017); Vishnubhotla et al. (2019); Šeļa et al. (2023).
With the recent introduction of PDNC Vishnubhotla et al. (2022), a corpus of 28 novels annotated with quotation attribution, analysis on fictional characters in novels can be done at a larger scale. Michel et al. (2024) provide a comparative study of PDNC’s characters, converting the original PDNC quotation attribution task to a CV task and showing that UAR Rivera-Soto et al. (2021) performs well on the task. However, this model is trained on Reddit posts, which are inherently different from fictional character utterances, motivating the need for an in-domain variant of UAR.
text
Quotation Attribution.   Datasets of quotation attribution containing fully annotated English novels are rather scarce, and are often limited to one or two books. Elson and McKeown (2010) introduce the CQSC corpus and attribute automatically extracted quotes to named entities and nominals with a supervised mention ranking model. Instead, He et al. (2013) attribute quotes directly to speakers with a supervised ranking system using features such as speaker alternation patterns and character-level features He et al. (2010). Other approaches involve sequence labeling O’Keefe et al. (2012) and Dialogue State Tracking Cuesta-Lazaro et al. (2022). The deterministic sieve-based model of Muzny et al. (2017b) regards quotation attribution as a two-step process: quote-mention linking and mention-speaker linking. The NLP pipeline dedicated to books, BookNLP, replaced the deterministic sieves with fine-tuned language models. The link between mention and speaker is carried out by BookNLP’s coreference resolution model, and the quote-mention linking is predicted by a BERT model, fine-tuned on LitBank’s speaker annotations Sims and Bamman (2020). Recently, Vishnubhotla et al. (2022) introduce the largest-to-date corpus of quotation attribution, PDNC, and show a similar accuracy score of around 63% for both BookNLP and the sieve-based model. However, better results were obtained later by fine-tuning BookNLP on PDNC and restricting the predicted coreference chains to resolved characters only Vishnubhotla et al. (2023). SIG Su et al. (2023) adapted quotation attribution for encoder-decoder models, showing that predicting both the speaker and the addressees of a quote with BART reaches state-of-the-art accuracy on PDNC.
@@
3 Literary Character Representations
In this section, we introduce DramaCV (Drama Character Verification), a dataset of English drama plays focused on CV. Then, we compare variants of Universal Authorship Representation (UAR) trained on DramaCV with the original UAR and SBERT. We finish by analyzing domain adaptation results on the CV variant of PDNC.
@@ 3 Literary Character Representations
Table 1: Summary statistics of the Scene and Play instances of DramaCV by split and the Explicit split of Michel et al. (2024) that we use to test generalization.
Among models trained on Authorship Verification, UAR, a fine-tuned variant of SBERT, is able to distinguish millions of Reddit authors. It learns a representation space under which a collection of documents composed by an author has a higher similarity with another collection from the same author than with a collection composed by another author Rivera-Soto et al. (2021). Although the Reddit version of UAR appears to perform well on PDNC, which is on-par with its zero-shot domain transfer abilities Rivera-Soto et al. (2021), we question its suitability for literary analysis and propose to train instead a domain-specific variant of UAR tailored for literary texts. Reddit has the advantage of spanning millions of different authors, but posts are often written informally and exhibit strong stylistic cues such as typographical or case errors Wegmann et al. (2022). In contrast, literary texts do not contain such errors and often span a wider vocabulary that evolves over time, in concordance with writing styles. This motivates the need for an in-domain UAR for analyzing authorship of fictional characters. Besides, quotes are subject to the writing style of authors. We would like character representations disentangled from the author’s stylistic preferences.
@@ 3.1 DramaCV
Character Verification aims at answering the following problem: given two collections of quotes that have been uttered by characters within the same work, predict whether they were spoken by the same character or not. Although closely linked to Authorship Attribution, we note that this framework is different for the following reasons: 1) we do not have access to a list of canonical characters for every literary work and 2) we compare characters within the same story, implying that each quote is inherently influenced by the author’s style. In comparison, Authorship Attribution aims at determining who wrote a text among a list of predefined authors, which is also inherently different from quotation attribution. Multiple documents come naturally annotated with who wrote or spoke them, such as social media Baumgartner et al. (2020) movie scripts Sang et al. (2022) or drama plays Fischer et al. (2019). Among existing datasets, DraCor Fischer et al. (2019) contains drama plays in various languages, but only provides Shakespeare plays for English language. Alternatively, explicit utterances from PDNC are an additional data source, but we wanted to evaluate the generalization of literary character embeddings for quotation attribution to unseen novels. Hence, we instead use the Explicit split of PDNC Michel et al. (2024) as an additional test split to analyze generalization from drama to novels. To the best of our knowledge, no datasets focusing on English drama and containing speaker annotation has been publicly released.
@@
We create DramaCV, a publicly available dataset of drama plays written by English authors, dating from the 15th to 20th century. We extract 499 plays from Project Gutenberg [^2] https://gutenberg.org/ with the GutenTag software Brooke et al. (2015) that we parsed to attribute each line to its speaker. Then, we construct two instances of DramaCV:
@@
Scene: We split each plays in scenes, a small segment unit of drama that is supposed to contain actions occurring at a specific time and place with the same characters. If a play has no <scene> tag, we instead split it in acts, with the <act> tag. Acts are larger segment units, composed of multiple scenes. We do not consider plays that have neither of these tags. Although scenes usually contain a small number of characters, they represent a challenge for CV models as participating characters are akin to discuss the same topic. We expect a model trained on scenes to rely mostly on stylistic cues.
Play: Here, we do not segment plays, and instead use all character lines in the full play. Compared to the scene instance, all characters participating in the play are considered as targets for the AV task. Here, topical preferences of each character across the whole play can be used in addition to style.
@@
We parsed a total of 169 plays for Scene, and 287 plays for Play. The resulting plays are split in 80% for training, 10% for validation and 10% for test. For validation and test, half utterances of a character are randomly sampled for the query, and the remaining half for the target. For the train split, we consider as queries and targets 8 utterances from the pool of a character’s set of lines, that we randomly sample at each epoch. We provide summary statistics for both instances of DramaCV in Table 1.
3.2 Experimental Setup
@@
We train UAR from scratch using its public implementation on DramaCV. For both instances, we use the pre-trained all-distilroberta-v1 as the base encoder, and employ a learning rate of 2e-5 for 20 epochs. We set the dimension of the model to $d=512$ and use a maximum sequence length of 64 tokens for all utterances. During training, we build queries and targets with distinct collections of 8 randomly sampled utterances to represent characters. A batch size of 1 play is used when training on the Play split, and 8 scenes are used when training on the Scene split. During inference, we do not restrict the number of utterances in a collection.
@@
Play. We want to avoid capturing the writing style of authors to distinguish among collections of character lines. Thus, we restrict the set of in-batch negatives to characters within the same work.
We compare our variants of UAR, $\text{UAR}_{Scene}$ and $\text{UAR}_{Play}$ against their Reddit counterpart, $\text{UAR}_{Reddit}$ and SBERT. Following previous work Aggazzotti et al. (2023); Michel et al. (2024), we create SBERT character embeddings by encoding each utterance in a vector with a pretrained model [^3] We use all-mpnet-base-v2., followed by a coordinate-wise mean of all utterance vectors. Instead, UAR directly encodes a collection of utterances into a $d$-dimensional vector.
For both test-splits of DramaCV, we encode a query with one of the mentioned models, yielding a character representation, and then encode each associated targets, which are collections of lines uttered by all characters in the same segment unit (including other lines from the same character). We then compute the cosine similarities between a query and the targets, and evaluate the ranking performance of these embeddings with the Area Under the Receiver Operating Characteristic curve (AUC) metric. We also test generalization to novels with the Explicit split of PDNC. In this experimental setup, Michel et al. (2024) proposed two evaluations: Character-Character (CC) and Character-Quote (CQ). The task in CC is to answer the following problem: given a collection of all explicit quotes of a character in a chapter, and a collection of quotes of any type from other chapters, determine if they have been spoken by the same character or not. The CQ experiment is constructed with the same collections of explicit quotes, but we instead need to predict if a single quote of any type from other chapters has been spoken by the same character or not. The later experiment resembles quotation attribution, but no contextual information is provided.
@@
Table 2: Average AUC (%) on test plays in DramaCV and novels in the Explicit split of PDNC with standard deviations between parentheses.
3.3 Results
We report AUC for both test splits of DramaCV and the Explicit split of PDNC in Table 2. Interestingly, $\text{UAR}_{Reddit}$ performs worse on both splits compared to SBERT and in-domain UARs. $\text{UAR}_{Scene}$ increases AUC on its training domain up to 10 points compared to SBERT, indicating that the model is able to capture stylistic cues specific to drama that help distinguishing utterances of characters discussing similar topics. The increase of $\text{UAR}_{Play}$ is slightly lower, but the high AUC indicates that it succeeds at distinguishing characters within the same play. The high standard deviation across plays suggests that task difficulty is not distributed uniformly among these literary works. Among best-performing plays, we found works by canonical authors such as Shakespeare’s Hamlet, Shaw’s John Bull’s Other Island or Ben Jonson’s The Alchemist (see complete list in Appendix A).
Looking at generalization results to PDNC, $\text{UAR}_{Reddit}$ performs better than SBERT, but almost equals $\text{UAR}_{Scene}$ in the CC setup. We find significant improvement with $\text{UAR}_{Play}$ over the baselines, suggesting that training UAR on drama allows better domain adaptation to literary novels. We also remark that this adaptation requires less data, as training $\text{UAR}_{Scene}$ with around 5000 drama characters is enough to match the performance of $\text{UAR}_{Reddit}$ that has been trained with millions of different users.
The advantages of domain adaptation can also be seen in the CQ evaluation setup, where the two variants of UAR trained on drama show a significant increase in AUC over the baselines. This indicates that character representations built with our proposed UARs are more similar to embeddings of their other quotes than representations built with $\text{UAR}_{Reddit}$ and SBERT. However, AUC scores for the CQ experiment are low, showing that the task of attributing quotes requires more than only comparing similarities of character and quote embeddings.
Motivated by the ability of UAR to encode character utterances in a representation space that contain useful, though insufficient information for quotation attribution, we evaluate next the impact of combining these global character embeddings with BookNLP’s contextual information.

### v2

@@ Abstract
Humans naturally attribute utterances of direct speech to their speaker in literary works. When attributing quotes, we process contextual information but also access mental representations of characters that we build and revise throughout the narrative. Recent methods to automatically attribute such utterances have explored simulating human logic with deterministic rules or learning new implicit rules with neural networks when processing contextual information. However, these systems inherently lack character representations, which often leads to errors in more challenging examples of attribution: anaphoric and implicit quotes. In this work, we propose to augment a popular quotation attribution system, BookNLP, with character embeddings that encode global stylistic information of characters derived from an off-the-shelf stylometric model, Universal Authorship Representation (UAR). We create DramaCV [^1] Code and data can be found at https://github.com/deezer/character_embeddings_qa, a corpus of English drama plays from the 15th to 20th century that we automatically annotate for Authorship Verification of fictional characters utterances, and release two versions of UAR trained on DramaCV, that are tailored for literary characters analysis. Then, through an extensive evaluation on 28 novels, we show that combining BookNLP’s contextual information with our proposed global character embeddings improves the identification of speakers for anaphoric and implicit quotes, reaching state-of-the-art performance.
@@ 1 Introduction
Recent approaches to quotation attribution typically succeed at attributing explicit quotes but struggle when it comes to anaphoric and implicit quotes Muzny et al. (2017b); Cuesta-Lazaro et al. (2022); Vishnubhotla et al. (2023); Su et al. (2023). Explicit utterances occur when the narrator indicates the speaker of a quote with a speech verb and a named mention, while anaphoric quotes are introduced with a speech verb and a pronoun or common noun. When no narrative information is given about the speaker of the quote, we refer to those as implicit quotes. Examples of such quotes are given in Figure 1.
@@
Attributing anaphoric and implicit quotes usually requires fine-grained understanding of contextual information such as discerning discussion patterns and long-range information such as linking pronouns to their canonical entity. Accessing information of fictional characters (style, persona, gender) also helps disambiguate challenging examples of attribution. Variations of writing style and long-range contextual dependencies, ubiquitous in fictional works, add an extra layer of complexity, making attribution of non-explicit quotes a complex task to solve even with recent tools. Sims and Bamman (2020); Vishnubhotla et al. (2022)
@@
Some of these shortcomings have been addressed by improving link prediction between pronouns/common nouns and canonical characters Vishnubhotla et al. (2023) or by exploiting better contextual representations Cuesta-Lazaro et al. (2022); Su et al. (2023). Instead, we show that combining fictional character representations encoding global stylistic information and topical preferences with contextual information processed by BookNLP,  [^2] https://github.com/booknlp/booknlp, a widely used NLP pipeline designed for literary texts, improves the speaker attribution of anaphoric and implicit quotes. As explicit utterances are often accurately attributed to their speaker Vishnubhotla et al. (2023), we assume those quotes given, and use them to derive stylistic information of characters.
@@
Among methods to represent characters from the words they utter, previous works Li et al. (2023); Aggazzotti et al. (2023); Michel et al. (2024) have explored embeddings encoding topical preferences using SBERT Reimers and Gurevych (2019) or stylistic information using Universal Authorship Representations (UAR) Rivera-Soto et al. (2021), a fine-tuned variant of SBERT trained on the Authorship Verification (AV) task with millions of Reddit users. UAR’s stylometric embeddings built from character utterances were previously shown to distinguish fictional characters among the same work Michel et al. (2024) but considering the discrepancies between Reddit and literary discourse, we question the effectiveness of these embeddings to represent characters. To overcome this potential limitation, we build variants of UAR tailored for literary texts that we validate on the AV task with drama and novel characters.
Stylometric character representations computed with our UAR variants significantly improve AV of characters in literary works, while also reaching state-of-the-art quotation attribution accuracy of non-explicit quotes when combined with BookNLP contextual representations. To sum up, our contributions are:
@@
  • We construct DramaCV, a corpus of English drama plays, dating from the 15th to 20th century, focused on the AV of fictional characters within the same work, and train variants of UAR on this dataset, showing that our proposed models significantly improve AV on DramaCV and also generalize better to novels.
@@
  • Through an extensive evaluation, we show that BookNLP can be easily extended to improve its current performance on the Project Dialogism Novel Corpus (PDNC) Vishnubhotla et al. (2022), and reaches state-of-the-art attribution accuracy of non-explicit quotes when trained in combination with global character representations computed from their explicit quotes.
@@ 2 Related Work
Quotation Attribution.
Datasets of quotation attribution containing fully annotated English novels are rather scarce, and are often limited to one or two books. Elson and McKeown (2010) introduce the CQSC corpus and attribute automatically extracted quotes to named entities and nominals with a supervised mention ranking model. Instead, He et al. (2013) attribute quotes directly to speakers with a supervised ranking system using features such as speaker alternation patterns and character-level features He et al. (2010). Other approaches involve sequence labeling O’Keefe et al. (2012), Dialogue State Tracking Cuesta-Lazaro et al. (2022) on English novels, or scoring systems on Chinese novels Chen2021. The deterministic sieve-based model of Muzny et al. (2017b) regards quotation attribution as a two-step process: quote-mention linking and mention-speaker linking. The NLP pipeline dedicated to books, BookNLP, replaced the deterministic sieves with fine-tuned language models. The link between mention and speaker is carried out by BookNLP’s coreference resolution model, and the quote-mention linking is predicted by a BERT model, fine-tuned on LitBank’s speaker annotations Sims and Bamman (2020). Recently, Vishnubhotla et al. (2022) introduce the largest-to-date corpus of quotation attribution, PDNC, and show a similar accuracy score of around 63% for both BookNLP and the sieve-based model. However, better results were obtained later by fine-tuning BookNLP on PDNC and restricting the predicted coreference chains to resolved characters only Vishnubhotla et al. (2023). SIG Su et al. (2023) adapted quotation attribution for encoder-decoder models, showing that predicting both the speaker and the addressees of a quote with BART reaches state-of-the-art accuracy on PDNC.
@@
Fictional Character Embeddings.
Understanding fictional characters objectives, personas and style often comes naturally when reading novels. When accessing information from the fictional world and interactions between characters provided throughout the narrative, we infer mental representations of characters that are revised when accessing new information Gernsbacher et al. (1992); Culpeper (1996). Previous computational works on this aspect of fictional character understanding have explored representing fictional characters in a latent space, often targeting a specific aspect, such as persona, style or descriptive elements. These computational representations of characters are often used for further analysis of large literary corpora, trying to understand character distinctiveness between authors Bamman et al. (2014) or narrative comprehension Brahman et al. (2021). Other works focus on using fictional characters embedings to improve character-level tasks such as coreference resolution and character linking Li et al. (2023) or contextual and book-level tasks such as narrative question answering Inoue et al. (2022).
Here, we propose to represent character from what they say and how they say it, building literary embeddings encoding style and topical preferences of characters, derived from models trained on the Authorship Verification (AV) of fictional characters within the same work. Most works exploring the AV of fictional characters focus on drama characters because of the availability of large annotated corpora and aim at capturing syntactic, lexical and phonological variations that occur in their direct speech using statistical measures of distinctiveness Dinu and Uban (2017); Vishnubhotla et al. (2019); Šeļa et al. (2023).
With the recent introduction of PDNC Vishnubhotla et al. (2022), a corpus of 22 novels annotated with quotation attribution3336 additional novels were released a year later, totaling 28 annotated novels., analysis on fictional characters in novels can be done at a larger scale. Michel et al. (2024) provide a comparative study of PDNC’s characters, converting the original PDNC quotation attribution task to the AV of fictional characters and showing that UAR Rivera-Soto et al. (2021) embeddings derived from a small collection of utterances are able to distinguish characters within the same work. However, Reddit – the target domain of UAR – is inherently different from traditional novels that use conventions on grammar, punctuation and format, motivating the need for a variant of UAR tailed for literary texts.
@@
Table 1: Summary statistics of the Scene and Play instances of DramaCV by split. A segment in the Scene split is a single scene or act in a play, while a segment in the Play split is the entire play. For each segment, the targets are collections of utterances uttered by all participating characters in the segment. The Explicit split of PDNC Michel et al. (2024) that we use to test generalization is also displayed for comparison.
3 Representing Fictional Characters
In this section, we present our approach to build and evaluate character representations tailored for literary texts, focusing on aspects of style and topical preferences of characters.
To build these representations, we train variants of UAR from scratch on the Authorship Verification (AV) of characters in the literary domain, considering a fictional character as an author in the AV terminology. The AV of fictional characters aims at answering the following problem: given two collections of quotes that have been uttered by characters within the same fictional work, predict whether they were spoken by the same character or not. Although closely linked to Authorship Attribution, we note that this framework is different for the following reasons: 1) we do not have access to a list of canonical characters for every literary work and 2) we compare characters within the same story, implying that each quote is inherently influenced by the author’s style.
UAR adapts SBERT by directly encoding collections of documents, and is trained to distinguish if two collections of documents have been written by the same author or not. Although the Reddit version of UAR appears to perform well on PDNC Michel et al. (2024), which is on-par with its zero-shot domain transfer abilities, we question its suitability for literary analysis and propose to train instead a domain-specific variant of UAR tailored for literary texts.
Only a few datasets containing annotations of the speaker of an utterance of direct speech in full novels exist, as annotating entire works is costly. Instead, prior works have considered using movie Azab2019; Sang et al. (2022); Li et al. (2023) or drama scripts Vishnubhotla et al. (2019); Fischer et al. (2019) as they come naturally annotated with the speaker of each utterance. We follow this approach and propose DramaCV (Drama Character Verification), a dataset of English drama plays focused on the AV of fictional characters. We chose to use drama plays as they are mostly similar to novels in terms of writing conventions, and because we were able to extract a number of characters and utterances an order of magnitude above the available datasets of annotated novels. The rest of this section evaluates our variants of UAR trained on DramaCV on the test set of DramaCV and on fictional characters participating in the 28 PDNC novels.
@@ 3.1 DramaCV
We extract 499 plays dating from the 15th to 20th century from Project Gutenberg [^4] https://gutenberg.org/ with the GutenTag software Brooke et al. (2015) that we automatically parsed to create a character list and to attribute each line to its speaker. Then, we construct two instances of DramaCV:
Scene.
@@
We split each play in scenes, a small segment unit of drama that is supposed to contain actions occurring at a specific time and place with the same characters. If a play has no <scene> tag, we instead split it in acts, with the <act> tag. Acts are larger segment units, composed of multiple scenes. For this split, we only consider plays that have at least one of these tags. Although scenes usually contain a few characters, they represent a challenge for CV models as participating characters are akin to discuss a similar topic. Thus, stylistic cues of characters are likely to be useful to distinguish characters within the same scene.
Play.
We do not segment play and use all character lines in a play. Compared to the scene segment, the number of candidate characters is higher, and discussions could include various topics. Hence, topical preferences of characters might be an additional useful feature to distinguish characters.
text
@@
After parsing, a total of 169 plays for the Scene split, and 287 plays for the Play split remained. For both splits, we use 80% of the plays for training, 10% for validation and 10% for test. Following the AV setup, we build queries and targets using collections of character utterances. In this setup, the objective is to recognize if a query and a target have been uttered by the same character or not. We construct queries in the validation and test splits by randomly sampling half of the utterances of a character, and we use the remaining half as the target. The train splits only consider characters that uttered at least 16 lines, such that we can build queries and targets with at least 8 utterances. More details on the construction of queries and targets can be found in Appendix A. Summary statistics for both instances of DramaCV are presented in Table 1.
3.2 Literary UAR
@@
We train UAR from scratch using its public implementation on DramaCV. For both instances, we use the pre-trained all-distilroberta-v1 as the base encoder, and employ a learning rate of 2e-5 for 20 epochs. We set the dimension of the model to $d=512$ and use a maximum sequence length of 64 tokens for all utterances. During training, we build queries and targets by randomly sampling 8 utterances to represent characters. A batch size of 1 play is used when training on the Play split, and 8 scenes are used when training on the Scene split. During inference, we do not restrict the number of utterances in a collection to reflect situations where we have access to only less than 8 utterances, or to more than 8 utterances.
@@
Play. We want to avoid capturing the writing style of authors to distinguish between collections of character lines. Thus, we restrict the set of in-batch negatives to characters within the same work.
We compare our variants of UAR, $\text{UAR}_{Scene}$ [^5] https://huggingface.co/gasmichel/UAR_Scene and $\text{UAR}_{Play}$ [^6] https://huggingface.co/gasmichel/UAR_Play against their Reddit counterpart, $\text{UAR}_{Reddit}$ and SBERT. Following previous work Aggazzotti et al. (2023); Michel et al. (2024), we create SBERT character embeddings by encoding independently each utterance in a feature $d$-dimensional vector [^7] We use all-mpnet-base-v2., followed by a coordinate-wise mean of all utterance vectors. Instead, UAR directly encodes a collection of utterances into a $d$-dimensional vector.
Given a collection of a character utterances (query) and collections of all characters occurring in the same segment (targets, that include a target uttered by the same character as the query), we evaluate the ability of these models to yield a higher cosine-similarity to the target uttered by the same character than to targets uttered by other characters using Area Under the Receiver Operating Characteristic curve (AUC). We use this evaluation setup on fictional characters for both test-sets of DramaCV and for the 28 novels of PDNC on the Explicit split of Michel et al. (2024). More details on the evaluation can be found in Appendix A.
@@
Table 2: Average AUC (%) on test plays in DramaCV and on novels in the Explicit split of PDNC with standard deviations between parentheses.
3.3 Authorship Verification Results
We report AUC for both test splits of DramaCV and the Explicit split of PDNC in Table 2. Interestingly, $\text{UAR}_{Reddit}$ performs worse on both splits compared to SBERT and in-domain UARs. $\text{UAR}_{Scene}$ increases AUC on its training domain up to 10 points compared to SBERT, indicating that the model is able to capture stylistic cues specific to drama that help at distinguishing utterances of characters discussing similar topics. The increase of $\text{UAR}_{Play}$ is slightly lower, but the high AUC indicates that it succeeds at distinguishing characters within the same play. The high standard deviation across plays suggests that task difficulty is not distributed uniformly among these literary works. Among best-performing plays, we found works by canonical authors such as Shakespeare’s Hamlet, Shaw’s John Bull’s Other Island or Ben Jonson’s The Alchemist (see complete list in Appendix B).
Looking at generalization results of PDNC, $\text{UAR}_{Reddit}$ performs better than SBERT, and almost equals $\text{UAR}_{Scene}$ in the Character-Character setup (CC), which compares similarity of character representations. We find significant improvement with $\text{UAR}_{Play}$ over the baselines, suggesting that training UAR on drama allows better domain adaptation to literary novels. We also remark that this adaptation is data-efficient, as training $\text{UAR}_{Scene}$ with around 5000 drama characters is enough to match the performance of $\text{UAR}_{Reddit}$ that has been trained on millions of different users.
The advantages of domain adaptation can also be seen in the Character-Quote (CQ) evaluation setup that aims at attributing utterances to their speaker by comparing the similarity between character and quote embeddings computed with the same model. Here, the two variants of UAR trained on drama show a significant increase in AUC over the baselines, indicating that character representations built with our proposed UARs are more similar to embeddings of their other quotes. However, AUC scores for the CQ experiment are low, showing that the task of attributing quotes requires more than only comparing similarities of character and quote embeddings.
Figure 2: Detailed process of BookNLP’s quotation attribution system. We inject extra-information in the form of character embeddings computed from explicit quotes $v_{\text{entity}(m)}$ in the fourth step, and quote embeddings $u_{q}$ in the second step.